OpenStudy (anonymous):

A gym offers regular memberships for $80 per month and off-peak memberships for $60 per month. Last month, the gym sold a total of 420 memberships for a total of $31,100. How many memberships sold were regular memberships? A)125 B)140 C)235 D)295

OpenStudy (anonymous):

@TrojanPoem

OpenStudy (trojanpoem):

Let regular memberships = y off-peak x x+ y = 420 80 y + 60x = 31100 You can have equations in two unknowns , solve for x , y and you 've the answer

OpenStudy (anonymous):

CAN YOU JUST HELP ME IM TIMED PLZZ

OpenStudy (anonymous):

I STILL HAVE 6 QUEESTIONS

OpenStudy (trojanpoem):

Well, the question says, it sold 420 membership ( we don't know which is regular and which is off-peak) but we know that off-peak membership sold + regular membership sold = 420 and the total gain is 31100 so 80 * number of off-peak sold + 60 * regular membership sold = 31100 (total money gained) for the ease I dubbed off-peak member ship y regular member ship x so i got x+y = 420 -> (1) 80y + 60x = 420 4y + 3x = 21 (2) Solve both equation (1), (2) x is your solution.

OpenStudy (anonymous):

THANK U CAN U HELP ME WITH MORE

OpenStudy (trojanpoem):

Sure.

OpenStudy (anonymous):

Solve the system of equations by substitution. What is the solution for x? 2x + y = 1 4x + 2y = −1 x = 0 x = 2 There is no x value, as there is no solution. x can be any value, as there is an infinite number of solutions.

OpenStudy (trojanpoem):

2x +y = 1 2x + y= -1 That doesn't make sense (2x +y) can't have two values at the same time so there is no x value as there is no solution to this system of equations again , you could use creamer rule to solve it or using matrices ranks as a better solution.

OpenStudy (anonymous):

WATS THE CREAMER RULE

OpenStudy (trojanpoem):